1: At a 10% discount rate, what’s the PRESENT value of a series of end-of-period cash flows equal to $700/period for 12 periods?

At a 10% discount rate, what’s the FINAL value of a series of end-of-period cash flows equal to $700/period for 12 periods?
Answer #1

a. Present worth = 700 * (P/A,10%,12) = 700 * 6.813692 = 4769.58 ~ 4770 (Nearest Dollar)

b. Future worth = 700 * (F/A,10%,12) = 700 * 21.384284 = 14968.998 ~ 14969 (Nearest Dollar)